OpenStudy (anonymous):

How would you rewrite the equation 3x+y=9 in y=mx+b form?

563blackghost (563blackghost):

https://www.youtube.com/watch?v=V6Xynlqc_tc this probably can help...

OpenStudy (texaschic101):

subtract 3x from both sides and what do you get ?

OpenStudy (photon336):

Step#1 isolate y by subtracting -3x from both sides 3x+y = y+3x 3x+y = 9-3x -3x now y = 9-3x